# ExoIris: fast exoplanet transmission spectroscopy in Python

### v1

@@ Abstract
I present ExoIris, a user-friendly Python package for exoplanet transmission and emission spectroscopy. Unlike existing tools, ExoIris models two-dimensional spectrophotometric transit time series directly and supports the joint analysis of multiple datasets obtained with different instruments and at different epochs. These features enable a self-consistent estimation of both wavelength-independent and wavelength-dependent parameters. They offer a more robust workflow compared to the commonly used two-step approach, where a ”white” light curve is fitted first and the transmission spectrum is then derived from independent fits constrained by the white-light solution. Despite its increased flexibility and robustness, ExoIris remains computationally efficient. A low-resolution transmission spectrum can be estimated from a single JWST NIRISS transit observation in $\sim 5$ minutes assuming white noise, and in $\sim 15$ minutes when accounting for time-correlated systematics using a Gaussian process noise model, on a standard desktop computer.
@@
software: PyTransit (Parviainen, 2015; Parviainen & Korth, 2020; Parviainen, 2020), LDTk (Parviainen & Aigrain, 2015), Astropy (Robitaille et al., 2013; Collaboration et al., 2018, 2022) emcee (Foreman-Mackey et al., 2013), celerite2 (Foreman-Mackey et al., 2023), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Numba (Lam et al., 2015).
@@ 1 Introduction
In ExoIris, wavelength-dependent quantities (planet-star radius ratio and stellar limb darkening profile) and wavelength-independent quantities (e.g., orbital elements) are inferred jointly from the full dataset. This contrasts with the common two-step approach, where each spectroscopic light curve is fitted independently, often using tight priors from a white-light curve analysis and fixed limb darkening coefficients from stellar atmosphere models. Such methods fail to use the information content of the observations fully and may underestimate uncertainties. By modeling the full two-dimensional dataset, ExoIris enables self-consistent inference of all relevant parameters, which results in more robust parameter estimates and transmission spectra.
ExoIris also supports joint modeling of heterogeneous datasets, allowing users to account for instrument- and time-dependent effects such as transit timing variations (TTVs), instrumental systematics (e.g. Moran et al., 2023; Alderson et al., 2025), and stellar contamination (e.g. Rackham et al., 2018, 2019; Rackham & de Wit, 2024). Time-correlated noise and systematics can be modeled with Gaussian processes (GPs), using fixed or inferred hyperparameters.
Despite its flexibility, ExoIris remains computationally efficient. The planet-star radius ratio is modeled as a spline over wavelength, with freely located knots treated as free parameters. This decouples the spectral resolution of the transmission spectrum from that of the data, enabling high-resolution sampling of sharp absorption features while allowing the smoother continuum to be modeled at lower resolution. The number of free parameters in the final model is primarily set by the number of radius ratio spline knots, and adding new datasets does not significantly increase model dimensionality if the wavelength space is already covered. For example, a joint fit to a single JWST NIRISS transit with $\mathcal{R}\sim 300$, 100 radius ratio knots, and a GP noise model typically completes in 10–20 minutes on a standard desktop computer. More complex analyses, such as the joint modeling of eight WASP-39 b Early Release Science datasets (Sect. 3), complete within a few hours.
2 ExoIris
2.1 Spectrophotometry Model Overview
@@
$F_{\mathrm{d}}(t,\lambda)=O+(1-O)B_{\mathrm{d}}(\lambda)\;T\left(t,k(\lambda),\mathbb{l}(\lambda),\boldsymbol{\Theta}\right),$ (1)
where $O$ is a flux offset term, $B$ is a baseline flux model, and $T$ is the spectrophotometric transit model. The model depends on the planet-star radius ratio $k$, the limb darkening parameters $\mathbb{l}$, and the orbital parameter vector $\boldsymbol{\Theta}$ (comprising the zero epoch $t_{0}$ for each epoch group, orbital period $p$, inclination $i$, eccentricity $e$, and argument of periastron $\omega$). Each dataset is assigned to offset, noise, epoch, and baseline model groups, which determine the model parameterization. These are described in detail in Sects. 2.5, 2.8, 2.7, and 2.6, respectively.
Figure 1: Example of a spectrophotometric transit model. The first two panels show the power-2 limb darkening coefficients as splines over wavelength with freely placed knots. The third panel shows the planet’s radius ratio modeled in the same way. The rightmost panel displays the model flux as a function of time and wavelength.
The planet-star radius ratio, $k$, is modeled as a spline over wavelength, decoupling the resolution of the transmission spectrum from the spectroscopic resolution of the data (Sect. 2.3). The limb darkening vector $\mathbb{l}$ depends on the chosen model. Currently, ExoIris supports two-parameter analytical limb darkening laws with coefficients modeled as splines and a numerical LDTkLDM-based model (Sect. 2.4). In the analytical case, $\mathbb{l}$ consists of limb darkening coefficients at the knot wavelengths; in the numerical case, it contains the stellar effective temperature, metallicity, and surface gravity.
Figure 1 illustrates a single spectrophotometric model for a dataset with 1000 exposures and 600 wavelength elements. The model flexibly captures both sharp and smooth variations in $k$, allowing high resolution where needed and lower resolution elsewhere.
2.2 Transmission Spectroscopy Transit Model
ExoIris uses the TSModel transit model from PyTransit v2.6 (Parviainen, 2015; Parviainen & Korth, 2020; Parviainen, 2020), specifically designed for transmission spectroscopy. It is based on the latest version of the RoadRunner model (Parviainen, 2020), which enables fast light curve evaluation with arbitrary radially symmetric stellar surface brightness profiles. This flexibility supports both analytical limb darkening laws and numerical profiles derived from stellar atmosphere models.
@@ 2.2 Transmission Spectroscopy Transit Model
Among these quantities, $I_{\star}$ and $A_{\mathrm{p}}$ are straightforward to calculate. The main computational challenge lies in evaluating $\hat{I}_{\mathrm{p}}$ efficiently for arbitrary limb darkening profiles. The RoadRunner model addresses this by expressing $\hat{I}_{\mathrm{p}}$ as a weighted sum over the stellar intensity profile, reducing the model evaluation to a dot product between two one-dimensional arrays multiplied by the area occulted by the planet:
@@
where $\mathbf{w}(k,b)$ is a weight vector dependent on the radius ratio $k$ and sky-projected planet-star center separation $b$ as described in Parviainen (2020), $\mathbf{l}$ is the discretised limb darkening profile, and $a(k,b)$ is the occulted area.
@@
Separating the limb darkening term from the geometric occultation enables high computational efficiency, making the RoadRunner model well suited for transmission spectroscopy. A transmission spectroscopy dataset derived from a spectroscopic time series consists of flux estimates in multiple narrow passbands, all sampled at identical points in time. This allows for several optimizations that significantly reduce the computational cost of the model:
@@
2.3 Radius Ratio Parameterization
@@
In transmission spectroscopy, the key quantity of interest is the wavelength-dependent planet-star radius ratio, $k(\lambda)$, which encodes the shape of the transmission spectrum. While the TSModel enables the efficient calculation of thousands of spectrophotometric transit light curves, treating the radius ratio in each passband as an independent free model parameter would result in exceedingly large-dimensional models.
@@
ExoIris addresses this by decoupling the data resolution from the spectral resolution of the retrieved transmission spectrum. As shown in Fig. 1, the radius ratio is modeled as a spline over wavelength, defined by a user-specified set of knots. The values at the knot positions are treated as free parameters, and the radius ratio for each passband is interpolated from the spline. Both the number and placement of the knots are fully customizable. In typical use, knots can be distributed uniformly or logarithmically in wavelength, or clustered around known or suspected spectral features. Although the transit model is evaluated at the full spectroscopic data resolution, the radius ratio values are constrained by the lower-dimensional spline.
@@
This parameterization offers several benefits. First, model complexity depends on the number of spline knots rather than the data resolution. Second, the spectral resolution of the transmission spectrum can be adapted to the science case—for example, by placing more knots near expected absorption features and fewer in smooth continuum regions. Finally, the knot configuration can be modified dynamically during analysis, allowing for iterative refinement: an initial coarse (low-resolution) spectrum can be followed by a higher-resolution fit using the earlier result as a starting point.
2.4 Limb Darkening
2.4.1 Analytical Limb Darkening Models
@@
ExoIris models the wavelength-dependent stellar limb darkening coefficients using the same spline-based approach as for the planet-star radius ratio: a user-defined set of wavelength knots defines a spline, and the coefficient values at these knots are treated as free parameters (see Fig. 1). The limb darkening coefficients for individual passbands are then obtained by interpolation.
@@
This parameterization is available for all two-parameter analytical limb darkening laws supported by the RoadRunner model, including the quadratic and power-2 laws. It provides the same flexibility and regularization benefits as the spline-based radius ratio model: the number and placement of knots can be adjusted to balance spectral resolution against model complexity, and the knot configuration can be updated at any point during the analysis.
@@
As with the radius ratio, users can increase the knot density in regions where limb darkening is expected to change rapidly, or reduce the number of free parameters in smoother spectral regions. When applied to a joint fit of all spectrophotometric light curves, this approach reduces reliance on external priors and allows limb darkening constraints to be derived empirically from the data. Although ExoIris supports the use of priors generated with the LDTk package (Parviainen & Aigrain, 2015)–which uses PHOENIX stellar atmosphere models (Husser et al., 2013) to generate theoretical limb darkening profiles–these priors can be made broad enough to ensure that the inference is driven by the data, not the theory, when discrepancies arise.
2.4.2 LDTk-Based Limb Darkening Model
@@
In addition to the spline-based analytical models, ExoIris supports a numerical limb darkening model based on LDTk. This method uses PHOENIX stellar atmosphere models (Husser et al., 2013) to generate intensity profiles directly from three stellar parameters: effective temperature ($T_{\mathrm{eff}}$), surface gravity ($\log g$), and metallicity ($\mathrm{Fe}/\mathrm{H}]$). These parameters are treated as global free parameters, independent of the number of passbands, providing a low-dimensional and physically motivated alternative to spline interpolation.
@@
While the LDTk-based model is more tightly constrained by stellar theory, the spline-based approach offers greater flexibility when empirical, data-driven constraints are preferred. Both methods are fully supported in ExoIris and can be chosen based on the desired balance between theoretical and observational constraints.
2.5 Additive Flux Offset
@@
The offset term, $O$, in Eq. (1) accounts for additive bias-level variations arising from instrumental or astrophysical sources. These variations can be instrument- or epoch-dependent and directly affect the measured transit depths.
First, bias estimation errors in the instrument can lead to apparent depth differences between detectors, as seen in the NIRSpec observations of TOI-776 b (Moran et al., 2023; Alderson et al., 2025), where nonphysical jumps were observed between the transmission spectra from the NRS1 and NRS2 detectors. Second, the presence of unocculted stellar inhomogeneities, such as faculae or star spots, can cause wavelength-independent, epoch-dependent variations in the observed flux. This so-called ”transit light source effect” (Rackham et al., 2018, 2019; Rackham & de Wit, 2024) can imprint spectroscopic features that mimic atmospheric signals. While the effect is inherently wavelength-dependent, its relative variation across epochs can be approximated (to first order) as a constant offset in flux. Third, variable third-light contamination from an unresolved background source (e.g., a nearby variable star) can similarly produce epoch-dependent flux offsets.
@@
ExoIris addresses these effects using an offset parameter, $O$, which modifies the light curve model in Eq. (1). Each dataset is assigned to an offset group, and a separate $O_{i}$ parameter is introduced for each group. Offset group 0 is the reference and has its offset fixed to zero, while the offsets for other groups are treated as free parameters. The offset $O$ affects the overall transit depth uniformly across all wavelengths within a group, leaving the relative shape of the transmission spectrum unchanged.
@@
This implementation allows the model to absorb detector- or epoch-specific flux shifts without distorting the spectroscopic features of interest. It also provides a useful diagnostic: non-zero offsets may indicate instrumental or astrophysical systematics at the group level. However, the method relies on sufficient wavelength overlap between offset groups—either directly with the reference group or indirectly through overlapping intermediate groups. Without such overlap, the offset becomes degenerate with the radius ratio, and additional constraints are needed to resolve the degeneracy.
@@
Figure 2 illustrates this functionality using a simplified version of the TOI-776 b dataset from Alderson et al. (2025), where two transits were observed with both the NRS1 and NRS2 detectors. In the second transit, a shift is visible between the NRS1 and NRS2 light curves, while no such shift appears in the first transit. Datasets 0 and 1 correspond to the first epoch and are assigned to the reference offset group (group 0). Datasets 2 and 3, from the second epoch, are assigned to offset groups 1 and 2, respectively, introducing two additional offset parameters. In the example shown, the model is evaluated with offset values of 0.0 and 0.5 for groups 1 and 2, resulting in identical flux models for datasets 0 and 2. However, the apparent transit depths for dataset 3 are reduced by 50% relative to dataset 1, even though the physical radius ratio remains unchanged.
@@
Figure 2: Example illustrating the use of offset groups to account for instrument- or epoch-dependent bias-level variations arising from bias estimation errors or unocculted stellar inhomogeneities. The figure shows four spectrophotometric datasets from two transit epochs, each observed with two detectors. The first two columns show the spectrophotometric model for all datasets, the next two columns display the average light curves, and the final column shows the radius ratio as a function of wavelength. Here, DS is the dataset index, BG the offset group index, and EG the epoch group index (see Sect. 2.7).
2.6 Multiplicative Flux Baseline
In addition to the additive offset term, the flux model includes a wavelength-dependent multiplicative baseline term, $B_{\mathrm{d}}(\lambda)$, which represents the out-of-transit flux level. As with the radius ratio and limb darkening coefficients, the baseline is modeled as a spline over wavelength. However, each dataset is assigned its own independent baseline model, with a customizable number of spline knots (see Fig. 3).
The baseline model is not intended to capture short-scale spectral features—though this can be approximated by increasing the number of knots—nor is it designed to account for time-dependent variability, which is more appropriately handled by the GP noise model (Sect. 2.8). Instead, its primary purpose is to propagate uncertainties in the baseline flux level into the inferred transmission spectrum, ensuring that these uncertainties are properly reflected in the final parameter estimates.
Figure 3: Example of a baseline flux model in an ExoIris analysis with two spectrophotometric datasets. Each dataset is associated with a multiplicative baseline model, represented as a spline over wavelength. The spline knot values are included as free parameters in the model.
2.7 Transit Timing Variations

### v2

@@ Abstract
I present ExoIris, a user-friendly Python package for exoplanet transmission and emission spectroscopy. Unlike existing tools, ExoIris models two-dimensional spectrophotometric transit time series directly and supports the joint analysis of multiple datasets obtained with different instruments and at different epochs, as well as modeling stellar spot crossings and the influence of unocculted heterogeneities (the transit light source effect). These features enable a self-consistent estimation of both wavelength-independent and wavelength-dependent parameters. They offer a more robust workflow than the commonly used two-step approach, in which a ”white” light curve is fitted first, and the transmission spectrum is then derived from independent fits constrained by the white-light solution. Despite its increased flexibility and robustness, ExoIris remains computationally efficient. A low-resolution transmission spectrum can be estimated from a single JWST NIRISS transit observation in $\sim 5$ minutes assuming white noise, and in $\sim 15$ minutes when accounting for time-correlated systematics using a Gaussian process noise model, on a standard desktop computer.
@@
software: PyTransit (Parviainen, 2015; Parviainen and Korth, 2020; Parviainen, 2020), LDTk (Parviainen and Aigrain, 2015), Astropy (Robitaille et al., 2013; Collaboration et al., 2018, 2022), emcee (Foreman-Mackey et al., 2013), celerite2 (Foreman-Mackey et al., 2023), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Numba (Lam et al., 2015).
@@ I Introduction
In ExoIris, wavelength-dependent quantities (planet-star radius ratio and stellar limb-darkening profile) and wavelength-independent quantities (e.g., orbital elements) are inferred jointly from the full dataset. This contrasts with the common two-step approach, in which each spectroscopic light curve is fitted independently, often using tight priors from white-light curve analysis and fixed limb-darkening coefficients from stellar atmosphere models. Such methods fail to fully exploit the information content of the observations and may underestimate uncertainties. By modeling the full two-dimensional dataset, ExoIris enables self-consistent inference of all relevant parameters, yielding more robust parameter estimates and transmission spectra.
ExoIris also supports joint modeling of heterogeneous datasets, allowing users to account for instrument- and time-dependent effects such as transit timing variations (TTVs) and instrumental systematics (e.g. Moran et al., 2023; Alderson et al., 2025). Time-correlated noise and systematics can be modeled with Gaussian processes (GPs), using fixed or inferred hyperparameters. The model can also include an arbitrary number of stellar spot crossings, as well as time-varying levels of contamination from unocculted spots and faculae (the transit light source effect, TLSE; Rackham et al., 2018, 2019; Rackham and de Wit, 2024).
Despite its flexibility, ExoIris remains computationally efficient. The planet-star radius ratio is modeled as a wavelength-dependent interpolating function defined by a set of freely placed knots, whose values are included as free parameters in the model, with several interpolation schemes available. This approach decouples the spectral resolution of the transmission spectrum from that of the data, enabling high-resolution sampling of sharp absorption features while allowing the smoother continuum to be modeled at lower resolution. The number of free parameters in the final model is primarily determined by the number of radius-ratio knots, and adding new datasets does not significantly increase the model’s dimensionality if the wavelength space is already covered. For example, a joint fit to a single JWST NIRISS transit with $\mathcal{R}\sim 300$, 100 radius-ratio knots, and a GP noise model typically completes in 10–20 minutes on a standard desktop computer. More complex analyses, such as the joint modeling of eight WASP-39 b Early Release Science datasets (Sect. III), complete within a few hours.
II ExoIris
II.1 Spectrophotometry Model Overview
@@
$F_{\mathrm{d}}(t,\lambda)=O_{\mathrm{d}}+(1-O_{\mathrm{d}})B_{\mathrm{d}}(t,\lambda)\;T\left(t,k(\lambda),\mathbf{l}(\lambda),\boldsymbol{\Theta}\right),$ (1)
where $O$ is a flux offset term, $B$ is a baseline flux model, and $T$ is the spectrophotometric transit model. The model depends on the planet-star radius ratio $k$, the limb-darkening parameters $\mathbf{l}$, and the orbital parameter vector $\boldsymbol{\Theta}$ (comprising the zero epoch $t_{0}$ for each epoch group, orbital period $p$, inclination $i$, eccentricity $e$, and argument of periastron $\omega$). Each dataset is assigned to offset, noise, epoch, and baseline model groups, which determine the model parameterization. These are described in detail in Sects. II.5, II.9, II.7, and II.6, respectively.
Figure 1: Example of a spectrophotometric transit model. The first two panels show the power-2 limb-darkening coefficients as splines over wavelength with freely placed knots. The third panel shows the planet’s radius ratio modeled in the same way. The rightmost panel displays the model flux as a function of time and wavelength.
The planet-star radius ratio, $k(\lambda)$, is modeled as a wavelength-dependent interpolating function, decoupling the resolution of the transmission spectrum from the spectroscopic resolution of the data (Sect. II.3). The limb-darkening vector $\mathbf{l}$ depends on the chosen model. Currently, ExoIris supports two-parameter analytical limb-darkening laws with coefficients modeled as splines and a numerical LDTkLDM-based model (Sect. II.4). In the analytical case, $\mathbf{l}$ consists of limb-darkening coefficients at the knot wavelengths; in the numerical case, it contains the stellar effective temperature, metallicity, and surface gravity.
Figure 1 illustrates a single spectrophotometric model for a dataset with 1000 exposures and 600 wavelength elements using cubic B-spline interpolation for radius ratio and limb-darkening coefficients. The model flexibly captures both sharp and smooth variations in $k$, allowing high resolution where needed and lower resolution elsewhere.
II.2 Transmission Spectroscopy Transit Model
ExoIris uses the TSModel transit model from PyTransit v2.6 (Parviainen, 2015; Parviainen and Korth, 2020; Parviainen, 2020), specifically designed for transmission spectroscopy. It is based on the latest version of the RoadRunner model (Parviainen, 2020), which enables fast light curve evaluation with arbitrary radially symmetric stellar surface brightness profiles. This flexibility supports both analytical limb-darkening laws and numerical profiles derived from stellar atmosphere models.
@@ II.2 Transmission Spectroscopy Transit Model
Among these quantities, $I_{\star}$ and $A_{\mathrm{p}}$ are straightforward to calculate. The main computational challenge lies in evaluating $\hat{I}_{\mathrm{p}}$ efficiently for arbitrary limb-darkening profiles. The RoadRunner model addresses this by expressing $\hat{I}_{\mathrm{p}}$ as a weighted sum over the stellar intensity profile, reducing the model evaluation to a dot product between two one-dimensional arrays multiplied by the area occulted by the planet:
@@
where $\mathbf{w}(k,b)$ is a weight vector dependent on the radius ratio $k$ and sky-projected planet-star center separation $b$ as described in Parviainen (2020), $\mathbf{l}$ is the discretized limb-darkening profile, and $a(k,b)$ is the occulted area.
@@
Separating the limb-darkening term from the geometric occultation enables high computational efficiency, making the RoadRunner model well suited for transmission spectroscopy. A transmission spectroscopy dataset derived from a spectroscopic time series consists of flux estimates in multiple narrow passbands, all sampled at identical points in time. This allows for several optimizations that significantly reduce the computational cost of the model:
@@
II.3 Radius Ratio Parameterization
@@
In transmission spectroscopy, the key quantity of interest is the wavelength-dependent planet-star radius ratio, $k(\lambda)$, which encodes the shape of the transmission spectrum. While the TSModel enables the efficient calculation of thousands of spectrophotometric transit light curves, treating the radius ratio in each passband as an independent free parameter would result in an exceedingly large-dimensional model.
@@
ExoIris addresses this by decoupling the data resolution from the spectral resolution of the retrieved transmission spectrum. As shown in Fig. 1, the radius ratio is modeled as an interpolating function over wavelength, defined by a user-specified set of knots. The values at the knot positions are treated as free parameters, and the radius ratio for each passband is interpolated using one of the implemented interpolation schemes. Both the number and placement of the knots are fully customizable, and the locations of any individual knots can also be included as free model parameters. In typical use, knots can be distributed uniformly or logarithmically in wavelength, or clustered around known or suspected spectral features. Although the transit model is evaluated at the full spectroscopic data resolution, the radius-ratio values are constrained by the lower-dimensional interpolating function.
@@
This parameterization offers several benefits. First, model complexity depends on the number of radius-ratio knots rather than on the data resolution. Second, the spectral resolution of the transmission spectrum can be tailored to the science case—for example, by placing more knots near expected absorption features and fewer in smooth continuum regions. Finally, the knot configuration can be modified dynamically during analysis, allowing for iterative refinement: an initial coarse (low-resolution) spectrum can be followed by a higher-resolution fit using the earlier result as a starting point.
Figure 2: Radius-ratio interpolation schemes supported by ExoIris. The black dots show the posterior median radius-ratio knots with their 68% central posterior intervals, as inferred from the JWST WASP-39 b NIRISS observations presented in Feinstein et al. (2023) using five of the supported interpolation schemes (shifted in y for visualization). The black lines show the posterior median transmission spectra, and the colored shading shows their 68% central posterior intervals.
Currently, ExoIris supports nearest-neighbor, linear, quadratic B-spline, cubic B-spline, PCHIP, and Makima interpolation (Fig. 2), all implemented via SciPy (Virtanen et al., 2020). The nearest-neighbor scheme models the transmission spectrum as a piecewise constant step function, effectively treating the regions around knots as discrete bins, while the linear scheme connects the knots with straight line segments. The linear option is generally the most robust choice for low-signal-to-noise (SNR) datasets. The quadratic and cubic B-spline schemes produce smooth spline curves, but they can introduce spurious oscillations (overshoot) in low-SNR regions or near sharp spectral features. To mitigate this while retaining smoothness, ExoIris offers the PCHIP and Makima options. These are shape-preserving, non-linear interpolators that enforce monotonicity to prevent artificial extrema, though they are computationally more expensive than the other schemes.
II.4 Limb Darkening
II.4.1 Analytical Limb-Darkening Models
@@
ExoIris models the wavelength-dependent stellar limb-darkening coefficients using the same interpolation-based approach as for the planet-star radius ratio: a user-defined set of knots defines an interpolating function, and the coefficient values at these knots are treated as free parameters (see Fig. 1).
@@
This parameterization is available for all two-parameter analytical limb-darkening laws supported by the RoadRunner model, including the quadratic and power-2 laws. It provides the same flexibility and regularization benefits as the interpolation-based radius-ratio model: the number and placement of knots can be adjusted to balance spectral resolution against model complexity, and the knot configuration can be updated at any point during the analysis.
@@
As with the radius ratio, users can increase the knot density in regions where limb darkening is expected to change rapidly, or reduce the number of free parameters in smoother spectral regions. When applied to a joint fit of all spectrophotometric light curves, this approach reduces reliance on external priors and allows limb-darkening constraints to be derived empirically from the data. Although ExoIris supports the use of priors generated with the LDTk package (Parviainen and Aigrain, 2015) (which uses PHOENIX stellar atmosphere models by Husser et al. 2013 to generate theoretical limb-darkening profiles), these priors can be made broad enough to ensure that the inference is driven by the data, not the theory, when discrepancies arise.
II.4.2 LDTk-Based Limb-Darkening Model
@@
In addition to the interpolation-based analytical models, ExoIris supports a numerical limb-darkening model based on LDTk. This method uses PHOENIX stellar atmosphere models (Husser et al., 2013) to generate intensity profiles directly from three stellar parameters: effective temperature ($T_{\mathrm{eff}}$), surface gravity ($\log g$), and metallicity ($\mathrm{Fe}/\mathrm{H}]$). These parameters are treated as global free parameters, independent of the number of passbands, providing a low-dimensional and physically motivated alternative to spline interpolation.
@@
While the LDTk-based model is more tightly constrained by stellar theory, the interpolation-based approach offers greater flexibility when empirical, data-driven constraints are preferred. Both methods are fully supported in ExoIris and can be chosen based on the desired balance between theoretical and observational constraints.
II.5 Additive Flux Offset
@@
Errors in data reduction, such as incorrect bias estimation or instrumental systematics, can lead to apparent depth differences between detectors, as suggested for the NIRSpec observations of TOI-776 b (Moran et al., 2023; Alderson et al., 2025), where non-physical jumps were observed between the transmission spectra from the NRS1 and NRS2 detectors. The offset term, $O$, in Eq. (1) can be used to account for additive bias-level variations arising from instrumental or astrophysical sources. In general, these variations can be instrument- or epoch-dependent and directly affect the measured transit depths.
@@
Each dataset is assigned to an offset group, and a separate $O_{i}$ parameter is introduced for each group. Offset group 0 serves as the reference and has its offset fixed to zero, while the offsets for other groups are treated as free parameters. The offset $O$ modifies the overall transit depth uniformly across all wavelengths within a group, leaving the relative shape of the transmission spectrum unchanged.
@@
This implementation allows the model to absorb detector- or epoch-specific flux shifts without distorting the spectroscopic features of interest. It also provides a useful diagnostic: non-zero offsets may indicate instrumental or astrophysical systematics at the group level. However, the method relies on sufficient wavelength overlap between offset groups, either directly with the reference group or indirectly through overlapping intermediate groups. Without such overlap, the offset becomes degenerate with the radius ratio, and additional constraints, such as informative priors on the offset parameters, are required to resolve the degeneracy. Alternatively, without wavelength overlap, the datasets can all be assigned to offset group 0 and the offset modeling can be carried out in the retrieval phase.
@@
Figure 3 illustrates this functionality using a scenario similar to that encountered by Alderson et al. (2025), where two transits were observed with both the NRS1 and NRS2 detectors. In the second transit, a shift is visible between the NRS1 and NRS2 light curves, while no such shift appears in the first transit. Datasets 0 and 1 correspond to the first epoch and are assigned to the reference offset group (group 0). Datasets 2 and 3, from the second epoch, are assigned to offset groups 1 and 2, respectively, introducing two additional offset parameters. In the example shown, the model is evaluated with offset values of 0.0 and 0.5 for groups 1 and 2, resulting in identical flux models for datasets 0 and 2. However, the apparent transit depths for dataset 3 are reduced by 50% relative to dataset 1, even though the physical radius ratio remains unchanged.
@@
Figure 3: Example illustrating the use of offset groups to account for instrument- or epoch-dependent additive offsets arising from, for example, errors in bias-level estimation or instrumental systematics. The figure shows four spectrophotometric datasets from two transit epochs, each observed with two detectors. The first column shows the radius ratio as a function of wavelength, the next two columns show the spectrophotometric model for all datasets, and the final two columns display the average light curves. The dataset index is denoted by DS, the offset-group index by BG, and the epoch-group index by EG (see Sect. II.7). Datasets 0 and 1 were observed at epoch 0 and do not feature an offset between the two detectors. However, datasets 2 and 3, observed at epoch 1, show an offset, with the transit of dataset 3 being, on average, half the transit depth of dataset 1.
II.6 Multiplicative Flux Baseline
In addition to the additive offset term, the flux model includes a wavelength-dependent multiplicative baseline term, $B_{\mathrm{d}}(t,\lambda)$, which accounts for both the out-of-transit flux level and deterministic systematics. The baseline is modeled as a linear combination of dataset-specific covariates that, by default, include a constant intercept and a user-defined number of polynomial time terms centered on the dataset’s mid-time. Furthermore, the user can incorporate any relevant covariates for each dataset as required. For example, common-mode systematics estimated from an initial white-light-curve fit (a common practice in ground-based transmission spectroscopy), or auxiliary variables observed simultaneously with the photometry.
The baseline coefficients are determined through analytic profiling: during each log-posterior evaluation, a linear least-squares fit is performed against the multiplicative residuals (defined as the observed flux divided by the current transit model) to identify the best-fit baseline for the proposed transit parameters. By analytically resolving the linear baseline coefficients, ExoIris ensures that baseline trends are self-consistently accounted for (and profiled out during posterior sampling) without increasing model dimensionality.
II.7 Transit Timing Variations